OpenStudy (anonymous):

find the sum of natural numbers 1+2+3+4..110 up to 110

OpenStudy (anonymous):

\[\frac{110\times 111}{2}\]

OpenStudy (anonymous):

general formula is \[\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

6105